Need help with the following:
For the given functions f and g, find the requested function and state its domain.
f(x) = sqrt x + 11; g(x) = 3/x
Find f ∙ g.

Answer by Theo(13342)

f o g = f(g(x)) which means f of (g of x).

g(x) = x/3

f(x) = sqrt(x) + 11

f(g(x)) answers the question "what is the value of the equation f(x) = sqrt(x) + 11 if we replace x with g(x)?

so we do that.

we get f(g(x)) = sqrt(g(x)) + 11

since we know that g(x)) = 3/x, then we can replace g(x) with 3/x to get:

f(g(x)) = sqrt(3/x) + 11.

that's our answer.